OpenStudy (anonymous):

Okay, so I have an idea of where to start here, but it's finding the sine values that are bothering me. The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. C = 38°, a = 19, c = 10

OpenStudy (rhystic):

what exactly about the sine values?

OpenStudy (anonymous):

you are calculating side C?

OpenStudy (anonymous):

oh nvm B

OpenStudy (anonymous):

?

OpenStudy (anonymous):

When solving the equation, they'd like me to look for all values of ABC, as well as side a.

OpenStudy (anonymous):

I dont think a triangle can be formed

OpenStudy (anonymous):

so what side are we trying to calculate though?

OpenStudy (anonymous):

Solving for sin c got me a decimal, as it always does. The issue here is that sin ^-1 (answer) "should" calculate the decimal for me. I'm attempting to calculate angles C, A, and side a.

OpenStudy (anonymous):

there is no triangle

OpenStudy (anonymous):

For retrice(SSA) theorem sin(C) > c/a so there are no solutions and no triangle!

OpenStudy (anonymous):

(retrice

OpenStudy (anonymous):

\[(retrice\]

OpenStudy (anonymous):

lol it wont let me put A S S

OpenStudy (anonymous):

¯\_(ツ)_/¯

OpenStudy (anonymous):

For A S S(SSA) theorem sin(c)>c/a so there are no solutions and no triangle

OpenStudy (rhystic):

Yeah Niguyver is correct.

OpenStudy (anonymous):

The way my lesson has instructed me to solve it, \[\sin b / b = \sin c /c \] which does give a decimal value, so I'm a little confused there.

OpenStudy (anonymous):

there is no B

OpenStudy (anonymous):

Wait, which one did I post... No, you're right. Crap. Got my questions mixed up.